# Global Constraint LLM Agents for Text-to-Model Translation

### v1

@@ Abstract
Natural language descriptions of optimization or satisfaction problems are challenging to translate into correct MiniZinc models, as this process demands both logical reasoning and constraint programming expertise. We introduce a framework that addresses this challenge with an agentic approach: multiple specialized large language model (LLM) agents decompose the modeling task by global constraint type. Each agent is dedicated to detecting and generating code for a specific class of global constraint, while a final assembler agent integrates these constraint snippets into a complete MiniZinc model. By dividing the problem into smaller, well-defined sub-tasks, each LLM handles a simpler reasoning challenge, potentially reducing overall complexity. We conduct initial experiments with several LLMs and show better performance against baselines such as one-shot prompting and chain-of-thought prompting. Finally, we outline a comprehensive roadmap for future work, highlighting potential enhancements and directions for improvement.
@@ 1 Introduction
Building correct MiniZinc models from natural language descriptions is a complex challenge. Recently, Singirikonda et al. (2025) has introduced datasets Text2Zinc for benchmarking this task. Baseline evaluations on Text2Zinc (using direct prompting, chain-of-thought, and compositional strategies) found that even powerful LLMs are “not yet a push-button technology” for generating combinatorial models from text. In other words, general-purpose prompting often fails to capture all variables and constraints correctly, especially for harder optimization problems. This motivates research into more structured and guided methods.
@@
However, in existing multi-agent approaches, each agent still inherits the full complexity of the problem rather than focusing on a narrower, more tractable sub-task. To address this, we propose a new agentic framework for translating text to MiniZinc (Nethercote et al., 2007) that is centered around global constraints. In constraint programming, global constraints like all_different or cumulative are high-level primitives that capture common patterns among variables. They offer expressive and efficient building blocks in models, and many optimization problems can be described as a combination of such global constraints. Our approach leverages this by dedicating a specialized LLM agent to each type of global constraint, turning model generation into a collaboration of focused experts rather than one monolithic generation. The sections below outline background and related work, our framework, preliminary findings, and future steps in this paper.
@@ 2 Background and Related Work
2.1 NL4OPT
LLMs are increasingly applied to optimization and constraint programming. Early work Ramamonjison et al. (2023) tackled linear programming via entity recognition and logical-form translation, showing promising ChatGPT results on NL4OPT. Holy Grail 2.0 (Tsouros et al., 2023) outlined a blueprint for conversational modeling assistants. Accuracy of LLM-generated MiniZinc improved with in-line entity annotations (Dakle et al., 2023). Agentic methods followed: a multi-agent Chain-of-Experts (Xiao et al., 2023) and Optimus (AhmadiTeshnizi et al., 2023), a modular system for complex descriptions. NL-to-constraint translation was also explored via simple decomposition prompts with GPT (Tsouros et al., 2023). Extending this, RAG-based in-context learning built CPMpy models (Michailidis et al., 2025). Concurrently to our work, Szeider (2025) introduce a Reason-and-Act framework solving all 101 CP-Bench tasks (Michailidis et al., 2025).
2.2 MiniZinc
@@ 2.3 Global Constraints
Global constraints concisely represent recurring patterns like all-differentness, resource limits, ordering, counting, etc., which are found across many scheduling, assignment, and configuration problems. For example, all_different enforces that a set of decision variables all have distinct values – a common requirement in scheduling and allocation problems. Other examples include cumulative (which enforces scheduling resource capacity over time) and global_cardinality (which limits how many variables take each value).
3 Method
@@
Specialized LLM Agents per Constraint: For each type of global constraint, we instantiate a separate LLM agent with a specialized prompt. This prompt primes the agent to act as an expert in detecting and formulating that constraint in MiniZinc. The agent receives the full problem description (and any input data) but its instructions are local: it should ignore the broader context and only answer the question, “Does this problem involve an <X> constraint? If yes, produce the MiniZinc snippet for it; if not, output FALSE.” Each agent is effectively performing a binary classification (constraint present or not) followed by code generation for that constraint if needed. Critically, the agent is told not to produce any other modeling elements beyond its constraint. For instance, the all_different agent is prompted: “You are a MiniZinc modeling assistant specialized in detecting and modeling all_different constraints. Given a problem description, decide whether it requires one or more all_different constraints. If it does, generate only MiniZinc code specifying the all_different constraint and its variables. If it does not, return FALSE with a brief reason.”. Similar templates are crafted for each global constraint type, incorporating definitions and common clue words (e.g. phrases like "each … different" hint at all_different, or "no overlap" hints at cumulative). By isolating each agent’s focus, we simplify its reasoning task — it does not need to understand the entire problem structure, only whether a specific pattern appears and how to encode it.
@@
Assembler LLM for Model Integration: Once the constraint-specific agents have each returned either a code snippet or FALSE, an assembler LLM takes over. The assembler’s input includes the original problem description and all the constraint snippets (we call them "hints") provided by the agents that found a constraint. The assembler’s role is to compile a complete, coherent MiniZinc model from these pieces. We prompt the assembler agent as if it were “the world’s best MiniZinc programmer” tasked with integrating hints and filling in the gaps. Concretely, it must: (1) declare all decision variables (and their domains) that are needed, possibly renaming or merging variables from different snippets for consistency; (2) analysis and decide if to include provided global constraints or not; (3) add any remaining constraints from the text that were not covered by the hints; (4) determine the objective (if an optimization problem) or a satisfy goal, based on the description; and (5) append a proper solve item and output format. The assembler may ignore an irrelevant hint if an agent was mistaken, but in general it tries to use all valid snippets. By design, this agent has the most complex task, since it sees the full problem and must ensure completeness. However, because much of the heavy lifting (expressing complex constraints) is done by the specialized snippets, the assembler can focus on glueing components together and writing boilerplate code.
@@
The overall workflow (Figure 1) is akin to a team of specialists: each constraint agent independently proposes part of the model, and the assembler (acting like a senior engineer) reviews and integrates these contributions into the final solution. This approach aims to reduce the cognitive load on any single LLM — each one works on a local modeling sub-problem instead of the entire NL-to-Model translation.
Figure 1: Agentic pipeline for global-constraint detection and assembler.
@@
Table 1: Detection and false-detection rates by global constraint type.
4 Preliminary Experiments
@@
We conducted an initial evaluation of the Global Constraint Agents framework, focusing on two aspects: (1) the ability of individual constraint-specific agents to correctly detect global constraints, and (2) the end-to-end performance of the agentic pipeline compared to baseline prompting strategies and chain-of-thought (CoT) (Wei et al., 2022).
4.1 Constraint Agent Detection
@@
Table 1 reports detection performance for seven global constraint types using the open-source LLM Phi4. We evaluate on all 567 instances from the Text2Zinc dataset (Singirikonda et al., 2025). For each constraint type, the detection rate is computed over the subset of instances whose ground-truth model contains that constraint. As shown in the table, the average detection rates is around 70% to 80%. To estimate the false detection rate, we use 110 verified Text2Zinc instances and exclude those whose ground-truth model includes the target constraint. Overall, our agents achieve strong detection rates, and false detection rates are generally low for rarer constraints; the main exception is count (28.3%), suggesting that distinguishing counting patterns from other numerical constraints remains challenging. These results indicate that the individual agents are reasonably effective, with room for improvement through stronger base models or improved prompting.
@@
4.2 End-to-End Pipeline Performance
We compare the global-constraint agentic system (agents) with direct prompting (baseline) and chain-of-thought (CoT) across three LLM configurations—OpenAI o3-mini (OpenAI, 2025b), gpt-4o-mini (OpenAI, 2023), and the open-source gpt-oss 20B model (OpenAI, 2025a) (Table 2). We executed each model and strategy on 110 verified instances from the Text2Zinc datasets (Singirikonda et al., 2025).
Agents consistently outperform CoT on the stronger models (o3-mini, GPT-4o-mini) across execution, solve, and mean score, and remain competitive on the 20B open model. While CoT gives a clear lift over the non-agent baseline on 20B, our first-pass global agents—with no model tuning, prompt optimization, or hyperparameter search—match or slightly exceed CoT on the stronger models and substantially improve executability overall. This supports the claim that the gains come from the decomposition/agentic assembly itself rather than prompt or parameter scaling (see Table 2).
@@ 5 Future Work
Optimize agents and output quality: Replace hand-crafted prompts with systematic optimization (automatic search, curated few-shot exemplars, adversarial negatives) and consider constraint-type fine-tuning to boost precision/recall. Pair these with compile-time snippet validation so each agent not only detects constraints correctly but also emits syntactically valid MiniZinc in isolation.
@@
Scale evaluation: Go beyond small models (e.g., Phi-4) by running stronger LLMs (e.g., GPT-4) and sweeping both open- and closed-weight models across sizes. Benchmark on datasets rich in global-constraint instances—currently about 70% of Text2Zinc lacks them—to better showcase the approach and reveal architectural headroom.
@@ 6 Conclusion
We introduced an agentic Text-to-Model pipeline that assigns specialized LLM agents to global constraints and uses an assembler to integrate their outputs. Even in a first pass—without tuning, prompt engineering, or hyperparameter search—our agents outperform a carefully curated CoT on stronger models and improve executability on the open 20B model, indicating that decomposition and agentic assembly are the key drivers of gains. The approach is modular and immediately extensible: finer handling of subtle constraints, compile-aware validation of snippets, and more robust staged assembly (conflict resolution, retry, and self-rectification) are straightforward next steps. With systematic error analysis and targeted feedback loops, we expect to turn this prototype into a dependable NL4OPT modeling strategy—and a strong default for text-to-optimization workflows.

### v2

# Gala: Global LLM Agents for Text-to-Model Translation

@@ Abstract
Natural language descriptions of optimization or satisfaction problems are challenging to translate into correct MiniZinc models, as this process demands both logical reasoning and constraint programming expertise. We introduce Gala, a framework that addresses this challenge with a global agentic approach: multiple specialized large language model (LLM) agents decompose the modeling task by global constraint type. Each agent is dedicated to detecting and generating code for a specific class of global constraint, while a final assembler agent integrates these constraint snippets into a complete MiniZinc model. By dividing the problem into smaller, well-defined sub-tasks, each LLM handles a simpler reasoning challenge, potentially reducing overall complexity. We conduct initial experiments with several LLMs and show better performance against baselines such as one-shot prompting and chain-of-thought prompting. Finally, we outline a comprehensive roadmap for future work, highlighting potential enhancements and directions for improvement.
@@ 1 Introduction
Building correct MiniZinc models from natural language descriptions is a complex challenge. Recently, Singirikonda et al. (2025) has introduced several modeling co-pilot frameworks as well as a supporting dataset, Text2Zinc, and corresponding leaderboard for benchmarking this task. Existing co-pilot evaluations on Text2Zinc (using direct prompting, chain-of-thought, and compositional strategies, and external tools such as Ner4Opt (Kadıoğlu et al., 2024), Knowledge Graphs) found that even powerful LLMs are “not yet a push-button technology” for generating combinatorial models from text. In other words, general-purpose prompting often fails to capture all variables and constraints correctly, especially for harder optimization problems. This motivates research into more structured and guided methods.
@@
The main drawback of existing multi-agent approaches is that each agent still inherits the full complexity of the problem rather than focusing on a narrower, more tractable sub-task. To address this, we propose a new agentic framework for translating text to MiniZinc (Nethercote et al., 2007) that is centered around global constraints. In Constraint Programming (CP), global constraints such as all_different and cumulative are high-level primitives that capture common patterns among variables. They offer expressive and efficient building blocks in models, and many optimization problems can be described as a combination of such global constraints.
Our approach, Gala: GlobAl LLM Agents leverages this by dedicating a specialized LLM agent to each type of global constraint, turning model generation into a collaboration of focused experts rather than one monolithic generation. More broadly, our approach can be viewed as aligning and combining the key strength of Constraint Programming (i.e., global constraints) with the Agentic Frameworks. The following sections outline background and related work, our framework, preliminary findings, and future roadmap.
@@ 2 Background and Related Work
2.1 Holy Grail 2.0, NL4Opt Ner4Opt, & Optimus
LLMs are increasingly applied to optimization and constraint programming. Holy Grail 2.0 (Tsouros et al., 2023) outlined a blueprint for conversational modeling assistants. Early work Ramamonjison et al. (2023) tackled linear programming via entity recognition and logical-form translation, showing promising ChatGPT results on NL4OPT.Ner4Opt showed that accuracy of LLM-generated MiniZinc improves with in-line entity annotations (Dakle et al., 2023) as also built into several co-pilot pipelines (Kadıoğlu et al., 2024). Agentic methods followed: a multi-agent Chain-of-Experts (Xiao et al., 2023) and Optimus (AhmadiTeshnizi et al., 2023), a modular system for complex descriptions. Text-to-model translation was also explored via simple decomposition prompts with GPT (Tsouros et al., 2023). Extending this, RAG-based in-context learning built CPMpy models (Michailidis et al., 2025). Concurrently to our work, Szeider (2025) introduce a Reason-and-Act framework solving all 101 CP-Bench tasks (Michailidis et al., 2025).
2.2 MiniZinc & Text2Zinc
@@ 2.3 CP & Global Constraints
In CP, global constraints concisely represent recurring patterns like all-differentness, resource limits, ordering, counting, etc., which are found across many scheduling, assignment, and configuration problems. For example, all_different enforces that a set of decision variables all have distinct values – a common requirement in scheduling and allocation problems. Other examples include cumulative (which enforces scheduling resource capacity over time) and global_cardinality (which limits how many variables take each value). Our approach Gala is designed to take advantage of CP global constraints in an agentic LLM architecture as detailed next.
3 Gala: Global LLM Agents
@@
Specialized LLM Agents per Constraint: For each type of global constraint, we instantiate a separate LLM agent with a specialized prompt. This prompt primes the agent to act as an expert in detecting and formulating that constraint in MiniZinc. The agent receives the full problem description (and any input data) but its instructions are local: it should ignore the broader context and only answer the question, “Does this problem involve an <X> constraint? If yes, produce the MiniZinc snippet for it; if not, output FALSE.” Each agent is effectively performing a binary classification (constraint present or not) followed by code generation for that constraint if needed.
Critically, the agent is instructed not to produce any other modeling elements beyond its constraint. For instance, the all_different agent is prompted: “You are a MiniZinc modeling assistant specialized in detecting and modeling all_different constraints. Given a problem description, decide whether it requires one or more all_different constraints. If it does, generate only MiniZinc code specifying the all_different constraint and its variables. If it does not, return FALSE with a brief reason.”. Similar templates are crafted for each global constraint type, incorporating definitions and common clue words (e.g. phrases like "each … different" hint at all_different, or "no overlap" hints at cumulative). By isolating each agent’s focus, our main novelty is to simplify the reasoning task leveraging CP global constraints. Unlike the previous agentic approaches, our agents do not need to understand the entire problem structure, only whether a specific pattern appears and how to encode it.
@@
Assembler LLM for Model Integration: Once the constraint-specific agents have each returned either a code snippet or FALSE, an assembler LLM takes over. The assembler’s input includes the original problem description and all the constraint snippets (we call them “hints”) provided by the agents that found a constraint. The assembler’s role is to compile a complete, coherent MiniZinc model from these pieces. We prompt the assembler agent as if it were “the world’s best MiniZinc programmer” tasked with integrating hints and filling in the gaps. Concretely, it must: (1) declare all decision variables (and their domains) that are needed, possibly renaming or merging variables from different snippets for consistency; (2) analysis and decide if to include provided global constraints or not; (3) add any remaining constraints from the text that were not covered by the hints; (4) determine the objective (if an optimization problem) or a satisfy goal, based on the description; and (5) append a proper solve item and output format. The assembler may ignore an irrelevant hint if an agent was mistaken, but in general, it tries to use all valid snippets. By design, this agent has the most complex task, since it sees the full problem and must ensure completeness. However, because much of the heavy lifting (expressing complex constraints) is done by the specialized snippets, the assembler can focus on glueing components together and writing boilerplate code.
@@
Figure 1 presents the overall architecture which is essentially a team of specialists: each constraint agent independently proposes part of the model, and the assembler (acting as the principal architect) reviews and integrates these contributions into the final solution. This approach aims to reduce the cognitive load on any single LLM. Eech agent works on a local model of the structured sub-problem, thanks to CP global constraints, instead of the entire text-to-model translation.
Figure 1: Gala: Global LLM Agents architecture for global-constraint detection and assembly.
@@
Table 1: Performance of Gala on detection and false-detection rates by global constraint type.
4 Initial Results
@@
We conduct an initial evaluation of Gala, focusing on two aspects: (1) the ability of global agents to correctly detect global constraints, and (2) the end-to-end performance of the agentic pipeline compared to baseline prompting strategies and chain-of-thought (CoT) (Wei et al., 2022).
4.1 Performance of Global Agents
@@
Table 1 reports detection performance for seven global constraint types using the open-source LLM Phi4. We evaluate on all 567 instances from the Text2Zinc dataset (Singirikonda et al., 2025). For each constraint type, the detection rate is computed over the subset of instances whose ground-truth model contains that constraint. As shown in the table, the average detection rates is around 70% to 80%. To estimate the false detection rate, we use 110 verified Text2Zinc instances and exclude those whose ground-truth model includes the target constraint.
As shown Table 1, overall, our agents achieve strong detection rates, and false detection rates are generally low for rarer constraints; the main exception is count (28.3%), suggesting that distinguishing counting patterns from other numerical constraints remains a challenge. These results indicate that the individual agents are reasonably effective, with room for improvement through stronger base models or prompt optimization.
@@
4.2 Performance of Gala on Text-to-Model Translation
In Table 2, we compare Gala with direct prompting (baseline) and chain-of-thought (CoT) across three LLM configurations: OpenAI o3-mini (OpenAI, 2025b), gpt-4o-mini (OpenAI, 2023), and the open-source gpt-oss 20B model (OpenAI, 2025a). We execute each model and strategy on 110 verified instances from the Text2Zinc datasets (Singirikonda et al., 2025).
Agents consistently outperform CoT on the stronger models (o3-mini, GPT-4o-mini) across execution, solve, and mean score, and remain competitive on the 20B open model. While CoT gives a clear lift over the non-agent baseline on 20B, our first-pass global agents, with no model tuning, prompt optimization, or hyperparameter search, match or slightly exceed CoT on the stronger models, and substantially improve execution success overall. This supports the claim that the gains come from the decomposition/agentic assembly itself rather than prompt or parameter scaling.
@@ 5 Future Work
Optimize Global Agents: Replace hand-crafted prompts with systematic optimization (automatic search, curated few-shot exemplars, adversarial negatives) and consider fine-tuning per global constraint to boost precision and recall. Pair these with compile-time snippet validation so each agent not only detects constraints correctly but also emits syntactically valid MiniZinc.
@@
Scale Evaluation: Go beyond small models (e.g., Phi-4) by running stronger LLMs (e.g., GPT-4) and sweeping both open- and closed-weight models across sizes. Benchmark on datasets rich in global-constraint instances, currently 70% of Text2Zinc lacks them, to better showcase the approach and reveal architectural headroom.
@@ 6 Conclusion
We introduce Gala: Global LLM Agents for Text-to-Model translation of satisfaction and optimization problems. We leverage specialized LLM agents to global constraints and an assembler to integrate them. Even in a first pass, without tuning, prompt engineering, or hyperparameter search, Gala outperforms a carefully curated CoT on stronger models and improve executability on the open 20B model, indicating that decomposition and agentic assembly are the key drivers of gains. The approach is modular and immediately extensible: finer handling of subtle constraints, compile-aware validation of snippets, and more robust staged assembly (conflict resolution, retry, and self-rectification) are natural next steps. With systematic error analysis and targeted feedback loops, we expect to turn this prototype into a robust and strong default for modeling co-pilots.